OPTICAL STARK AND PUMP/PROBE MICROWAVE OPTICAL DOUBLE RESONANCE SPEC-TROSCOPY OF GAS-PHASE CHROMIUM MONONITRIDE AND VANADIUM MONONITRIDE

\underline{J. SCOTT ROBINSON}, KEI-ICHI NAMIKI, D. M. GOODRIDGE and T. C. STEIMLE, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287.

High resolution optical spectroscopic studies of gas-phase chromium mononitride, CrN, and vanadium mononitride, VN, using molecular beam techniques have been performed. The R_{ee}(0.5) branch feature of the (0,0)A^{4}Π_{3/2} − X^{4}Σ_{1/2}^{−} band system for ^{52}CrN was recorded as a function of the static electric field in the range 1.2 - 2.0 kV/cm. The resultant Stark shifts were analyzed to produce permanent electric dipole moments of 2.31(4)D and 5.41(2)D for the ground X^{4}Σ_{1/2}^{−} and excited A^{4}Π_{3/2} states, respectively. The P_{e}(1), F′′ = 2.5 feature of the (0,0) D^{3}Π_{0} − X^{3}Δ_{1} band system for ^{51}V^{14}N (I=3.5) was also recorded as a function of static electric field in the range .4 - 1.2 kV/cm. The permanent electric dipole moments derived from a least squares analysis of the Stark shifts were 3.07(1)D for the ground X^{3}Δ_{1} state and 6.15(3)D for the excited D^{3}Π_{0} state. The ^{52}Cr^{14}N (I=1) hyperfine structure was determined from the analysis of 12 components of the lowest pure rotational levels using the pump/probe microwave-optical double resonance technique. The resulting parameters are (in cm^{−1}) B′′ = .62387360(74), B′ = .6060(1), γ′′ = .0070050(13), λ′′ = 2.611151(16), eqQ_{0}(^{14}N) = -.000025(10), b_{F} (^{14}N) = .0000062(34) and c(^{14}N) = -.000151(92). Comparisons to other experimental work and theoretical bonding models are given^{a}. Observed trends amongst the early transition metals will be discussed.

aWalter J. Balfour, Charles X. W. Qian and Chi Zhou, J. Chem. Phys. 106, 4383 (1997); James F. Harrison, J. Phys. Chem.

100, 3513 (1996); Margareta R. A. Blomberg and Per E. M. Siegbahn, Theor. Chim. Acta. 81, 365 (1992).